If A,B C are three ideals of a ring R then show that A(BC)=(AB)C


1
Expert's answer
2022-02-07T16:27:40-0500

Let us show that "A(BC)=(AB)C."

Let "x\in A(BC)." Then there exist "a\in A,\ b\in B,\ c\in C" such that "x=a(bc)." Since the mutiplicative operation of a ring is associative, we conclude that "x=a(bc)=(ab)c\in (AB)C." Therefore, "A(BC)\subset(AB)C."

On the other hand, let "x\in (AB)C." Then there exist "a\in A,\ b\in B,\ c\in C" such that "x=(ab)c." Since the mutiplicative operation of a ring is associative, we conclude that "x=(ab)c=a(bc)\in A(BC)." Therefore, "(AB)C\subset A(BC)."

It follows that "A(BC)=(AB)C."